# Runtime Analysis of the SMS-EMOA for Many-Objective Optimization

### 16 Dec 2023 v1

@@ 4.2 Runtime of SMS-EMOA
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with $\mu\geq M$ to optimize the $m$OJZJ problem. Then after at most $\frac{1}{2}em^{k+1}k^{k}$ iterations in expectation, the population (and also the populations afterwards) contains at least one inner Pareto optimum.
Proof.
This lemma is already proved if the initial population has at least one inner Pareto optimum. Now we consider that the initial population contains no inner Pareto optimum. Consider any population $P$ without inner Pareto optimum. Then we know that there exists a block $B_{i}:=[(i-1)n^{\prime}+1..in^{\prime}],i\in[1..m/2]$ such that for all $x\in P$, $A_{i}:=|x_{B_{i}}|_{1}\in[0..k-1]\cup[n^{\prime}-k+1..n^{\prime}]$. We regard it a success if $y$ with $|y_{B_{i}}|_{1}\in[k..n^{\prime}-k]$ is generated. Then the probability to generate such $y$ from $x$ with $A_{i}<k$ is at least
$\binom{n^{\prime}-A_{i}}{k-A_{i}}$ ${}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-(k-A_{i})}\mathopen{}\mathclose{{}\left(\frac{1}{n}}\right)^{k-A_{i}}\geq\frac{1}{e}\binom{n^{\prime}-A_{i}}{k-A_{i}}\mathopen{}\mathclose{{}\left(\frac{1}{n}}\right)^{k-A_{i}}\geq\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{n^{\prime}-A_{i}}{n(k-A_{i})}}\right)^{k-A_{i}}$
$\geq{}\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{n^{\prime}-k}{nk}}\right)^{k}\geq\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{n^{\prime}}{2nk}}\right)^{k}=\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{1}{mk}}\right)^{k},$
where the last inequality uses $k\leq n^{\prime}/2$ and the last equality uses $n^{\prime}=2m/n$. We also know that the probability to generate such $y$ from $x$ with $A_{i}>n^{\prime}-k$ is at least
$\binom{A_{i}}{A_{i}-(n^{\prime}-k)}$ ${}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-(A_{i}-(n^{\prime}-k))}\mathopen{}\mathclose{{}\left(\frac{1}{n}}\right)^{A_{i}-(n^{\prime}-k)}\geq\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{A_{i}}{n(A_{i}-(n^{\prime}-k))}}\right)^{A_{i}-(n^{\prime}-k)}$
$\geq{}\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{n^{\prime}-k}{nk}}\right)^{k}\geq\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{n^{\prime}}{2nk}}\right)^{k}=\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{1}{mk}}\right)^{k},$
where the last two inequalities use $n^{\prime}-k<A_{i}\leq n^{\prime}$ and $k\leq n^{\prime}/2$ respectively, and the last equality uses $n^{\prime}=2m/n$. Note that $A_{i}\in[0..k-1]\cup[n^{\prime}-k+1..n^{\prime}]$ holds for all $x\in P$ before such $y$ is generated. Then we know that $y$ will be generated in expected $e(mk)^{k}$ iterations. Since such $y$ cannot be dominated by any solution in $P$, we know that $y\in F_{1}$. From Lemma 4, we know that once $y$ is generated, $f(y)$ will be covered by the next population and by all future populations, and thus future populations will contain at least one solution $z$ with $|z_{B_{i}}|_{1}\in[k..n^{\prime}-k]$.
Since an inner Pareto optimum is reached when $y$ with $|y_{B_{1}}|_{1},\dots,|y_{B_{m/2}}|_{1}\in[k..n^{\prime}-k]$ is generated, we know that the expected iterations to reach an inner Pareto optimum is at most $\frac{m}{2}e(mk)^{k}=\frac{1}{2}em^{k+1}k^{k}$. ∎
We consider the next stage of covering all inner Pareto front points. The following lemma shows the runtime. The key of the proof is that one uncovered Pareto front point can be easily generated from a neighbor Pareto front if it is covered.
@@
Proof.
Let $u\in F^{*}$ be an uncovered inner Pareto front point that has some neighbor covered by the population $P$. Since such a neighbor is also a Pareto front point, we know that it is maintained in all future generations. Since the probability to generate an offspring with its function value of $u$ is at least $\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-1}\frac{1}{n\mu}\geq\frac{1}{en\mu}$, we know that the expected iterations to witness such an offspring is at most $n\mu$ in expectation. From Lemma 4, we know that once such an offspring is generated, $u$ will be covered in all future populations. Noting that there are at most $M$ inner Pareto front points, we know that in expected at most $en\mu M$ iterations, all Pareto front points are covered. ∎
The last stage is to cover the remaining Pareto front points. The following lemma shows the runtime. The key of the proof is that we can divide the Pareto front into several levels and any individual in the current $i$-th level can be easily generated from a point in the $(i-1)$-th level.
@@
Proof.
We call $(k,n^{\prime}+k)$ or $(n^{\prime}+k,k)$ extreme block. We divide the Pareto front points into $m/2+1$ levels, where each point in $i$-th level contains $i$ extreme blocks. Obviously, all inner Pareto front points belong to the 00-th level. Note that any point in $j$-th level ($j\geq 1$) can be generated by a certain point in $(j-1)$-th level via flipping specific $k$ bits with other bits unchanged. Hence, the probability to generate any point in $j$-th level condition on that all $(j-1)$-th level points are covered is at least
$\frac{1}{\mu}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-k}\mathopen{}\mathclose{{}\left(\frac{1}{n}}\right)^{k}\geq\frac{1}{e\mu n^{k}}.$
From Lemma 4, we know that such a Pareto front point is reached in at most $e\mu n^{k}$ in expectation, and will survive in all future populations. Consider the process of covering all non-inner Pareto front points in the order of covering the $1$-st level, …, $(m/2)$-th level. Since the 00-th level points are already covered from our assumption, and since there are at most $M$ non-inner Pareto front points, we know that the remaining levels are covered in at most $eM\mu n^{k}$ iterations in expectation. ∎
@@
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with $\mu\geq M$ to optimize the $m$OJZJ problem. Then after at most $\frac{1}{2}em^{k+1}k^{k}+n\mu M+eM\mu n^{k}$ iterations ($\mu+\frac{1}{2}em^{k+1}k^{k}+n\mu M+eM\mu n^{k}$ function evaluations) in expectation, the full Pareto front is covered.
Comparing with the NSGA-II’s exponential lower bound to cover the special case of $m$OJZJ, Theorem 8 shows the efficiency of the SMS-EMOA for many objectives, and indicates that for the NSGA-II, the non-dominated sorting is cool, but the crowding distance as the secondary criterion is deficient.
4.3 Runtime of GSMEO
For an overall comparison, we also show the runtime of the GSMEO on the $m$OJZJ. The GSEMO is a multiobjective counterpart of the $(1+1)$ EA. The initial population contains a uniformly and randomly generated solution. Then in each iteration, a solution is uniformly at random picked from the population to generate an offspring with standard bit-wise mutation. If this offspring cannot be dominated by any solution in the population, it will enter into the population and removes all solutions weakly dominated by it. It is not difficult to see all solutions in the population of the GSEMO are mutually non-dominated. Hence, the population size of the GSEMO is at most $M$ from Lemma 3.
The proof ideas are similar to Lemmas 5 to 7. The only difference is that SMS-EMOA needs Lemma 4 to ensure that all function values covered in $F_{1}$ will remain covered by the next population, but for GSEMO, all such function values are surely covered by the next population as $F_{1}$ contains all mutually non-dominated individuals whose function values will not be removed in the survival selection. Together with the upper bound $M$ of the population size, we have the following theorem.
@@ 4.3 Runtime of GSMEO
Let $k\leq n^{\prime}/2$. Consider using GSEMO to optimize the $m$OJZJ problem. Then after at most $\frac{1}{2}em^{k+1}k^{k}+nM^{2}+eM^{2}n^{k}$ iterations in expectation, the full Pareto front is covered.
5 Reduced Impact of Stochastic Population Updates
Bian et al. [BZLQ23] proposed a stochastic population update mechanism for the SMS-EMOA and proved that, somewhat unexpectedly given the state of the art, it can lead to significant performance gains. More precisely, they proved that the classic SMS-EMOA with reasonable population size $\mu=\Theta(n)$ solves the OJZJ problem in an expected number of $O(n^{k+2})\cap\Omega(n^{k+1})$ function evaluations, whereas for the SMS-EMOA with stochastic population update, $O(n^{k+2}\min\{1,2^{-k/4}n\})$ suffice. Hence, for $k=\omega(\log n)$, a super-polynomial speed-up is shown.
In this section, we extend the analysis of [BZLQ23] to $m$ objectives. Unfortunately, we will observe that the larger population sizes necessary here reduce the impact of the stochastic population update. As in [BZLQ23], we have no proven lower bounds for the SMS-EMOA with stochastic population update, but our upper-bound proofs suggest that the reduced effect of the stochastic population update on the runtime guarantee is real, that is, the impact on the true runtime is diminishing with the larger population sizes necessary in the many-objective setting.
5.1 Stochastic Population Update
The rough idea of the stochastic selection proposed by Bian et al. [BZLQ23] is that a random half of the individuals survive into the next generation regardless of their quality. The individual to be discarded is chosen from the remaining individuals according to non-dominated sorting and hypervolume contribution as in the classic SMS-EMOA. This approach resembles the random mixing of two acceptance operators in the Move Acceptance Hyper-Heuristic studied recently [LOW23, DDLS23].
More precisely, after generating the offspring $x^{\prime}$, the SMS-EMOA with stochastic selection chooses from $P_{t}\cup\{x^{\prime}\}$ a set $R^{\prime}$ of $\lfloor(\mu+1)/2\rfloor$ solutions randomly with replacement. The individual $z^{\prime}$ to be removed from $P_{t}\cup\{x^{\prime}\}$ is then determined via non-dominated sorting and hypervolume contribution applied to $R^{\prime}$ only. We note that with this mechanism, even the worst solution can enter the next generation with probability at least $1/2$.
5.2 Runtime
To establish the runtime result of the SMS-EMOA with stochastic population update, we separate the key facts into several lemmas for a clear demonstration. The following two lemmas 10 and 11 respectively show the general survival probability of each individual in the combined population, and the survival of the individual with rank one.
@@ 5.2 Runtime
Consider using SMS-EMOA with stochastic population update to optimize the $m$OJZJ problem. For any $x\in R_{t}$ for the current generation $t$, we have that there is $y$ with $f(y)=f(x)$ in $P_{t+1}$ with probability at least $1/2$.
Proof.
We simply calculate the probability of $x$ being removed after the population update. Since $x$ is possible to be removed only if it is selected in $R^{\prime}$, we know that it will be removed with probability at most $\lfloor|R^{\prime}|/2\rfloor/|R^{\prime}|\leq 1/2$, and hence $f(x)$ will be maintained in $P_{t+1}$ with probability at least $1/2$. ∎
@@
Consider using SMS-EMOA with stochastic population update and with $\mu\geq 2(M+1)$ to optimize the $m$OJZJ problem. For any $x\in F_{1}$ for the current generation $t$, we have that there is $y$ with $f(y)=f(x)$ in $P_{t+1}$. Specifically, any Pareto front point will be maintained in all future populations once it is reached.
Proof.
Let $x\in F_{1}$. If $\cup_{i\geq 2}F_{i}\neq\emptyset$, then the removal will happen in $\cup_{i\geq 2}F_{i}$, and $x\in P^{t+1}$.
If all solutions are in $F_{1}$, from Lemma 3, we know that $|f(F_{1})|\leq M$. Since there are $|R^{\prime}|\geq\lfloor(\mu+1)/2\rfloor\geq M+1>M$, we know that there exist $z_{1},z_{2}\in F_{1}$ such that $f(z_{1})=f(z_{2})$. Let $S_{x}=\{z\in R^{\prime}_{t}\mid f(x)=f(z)\}$. Then $|S_{x}|\geq 1$. If $|S_{x}|=1$, then $\Delta(x,F_{1})>0=\Delta(z_{1},F_{1})$, and thus $x\in P_{t+1}$. If $|S_{x}|>1$, since the survival selection only removes one solution, then there exists $y\in S_{x}$, such that $y\in P_{t+1}$.
Since any Pareto optimum is always in $F_{1}$, the latter part of this lemma holds. ∎
Recalling the proof in Lemmas 5 and 6, we know that these lemmas hold merely relies on that $f(F_{1})$ value will be maintained once it is reached. Hence, from Lemma 11, we have the following results for obtaining a population containing at least one inner Pareto optimum, and for covering all inner Pareto front points starting from a population with at least one inner Pareto optimum.
@@
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with stochastic population update and with $\mu\geq 2(M+1)$ to optimize the $m$OJZJ problem. Then
  • after at most $\frac{1}{2}em^{k+1}k^{k}$ iterations in expectation, the population (and also the populations afterwards) contains at least one inner Pareto optimum;
  • after at most $en\mu M$ iterations in expectation, all inner Pareto front points are covered starting from a population with at least one inner Pareto optimum.
Now we consider the runtime for the full coverage of the Pareto front after all inner Pareto front points are covered. Similar to the proof of Lemma 7 for the original SMS-EMOA, we partition the Pareto front points into $m/2+1$ levels based on the number of extreme blocks inside. The difference here is that we will consider the effect of the gap points, similar to [BZLQ23].
@@
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with stochastic population update and with $\mu\geq 2(M+1)$ to optimize the $m$OJZJ problem. Assume that the current population covers all inner Pareto front points. Then after at most $O(\min\{\frac{\mu k^{1/2}}{2^{k/2}},1\}M\mu n^{k})$ iterations in expectation, the full Pareto front is covered.
Proof.
We use the same definitions of the $j$-th ($j\in[0..m/2]$) level where each Pareto front point contains $j$ extreme blocks, and note that 00-th level points are covered. Let $P_{t}$ be the current population. We consider generating any point in $j$-th level condition on that all $(j-1)$-th level points are covered. Let $v$ be an uncovered Pareto front point in the $j$-th level. Then there exists a block $B$ that $v$ has the value of $(k,n^{\prime}+k)$ or $(n^{\prime}+k,k)$. W.l.o.g., let the block value of $v$ is $v_{B}=(k,n^{\prime}+k)$. Besides, there exists a set of points $U$ in the $(j-1)$-th level such that each $u\in U$ is the same as $v$ except that $u$ has the value of $(a,n^{\prime}+2k-a)$ with $a\in[2k..n^{\prime}]$ in the above block $B$. Let $S_{v}=\{f(z)\mid z\in P,z\text{\leavevmode\nobreak\ is the same as\leavevmode\nobreak\ }v\text{\leavevmode\nobreak\ except\leavevmode\nobreak\ }z_{B}\neq(k,n^{\prime}+k)\}$. Obviously, $y:=\min_{z\in S_{v}}z_{B}\leq 2k$. Let $A_{1}$ denote the event $k<y\leq\frac{3}{2}k$, $A_{2}$ be the event $y\geq\frac{7}{4}k$, and $A_{3}$ be $\frac{3}{2}k<y<\frac{7}{4}k$. Consider the potential function
$h(P_{t})=\begin{cases}0,&{}\text{if $v\in P_{t}$},\\ e\mu n^{k/2},&{}\text{if $A_{1}$},\\ e\mu n^{k/2}+1,&{}\text{if $A_{2}\cup A_{3}$}.\end{cases}$
For $A_{1}$, we pessimistically consider the decrease of $h$ by generating $v$ and the increase of $h$ by kicking out a corresponding solution of $y$. Hence, we have
$E[h$ ${}(P_{t})-h(P_{t+1})\mid P_{t},A_{1}]\geq\frac{1}{\mu}\frac{1}{n^{y-k}}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-(y-k)}e\mu n^{k/2}-\frac{1}{2}\cdot 1\geq 1-\frac{1}{2}=\frac{1}{2}.$
For $A_{2}$, $h$ will not increase, and there are $y-k\in[\frac{3}{4}k,k]$ ones in the block of a corresponding solution of $y$. We pessimistically consider the decrease of $h$ by flipping $y-k-\frac{1}{2}k$ ones and keeping other bits unchanged of the corresponding solution. Hence, we have
$E[h$ ${}(P_{t})-h(P_{t+1})\mid P_{t},A_{2}]\geq\frac{1}{\mu}\frac{\binom{y-k}{y-k-k/2}}{n^{y-k-k/2}}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-(y-k-k/2)}\frac{1}{2}$
[…]
where the last inequality uses $\frac{3}{4}k\leq y-k\leq k$. For $A_{3}$, $h$ will not increase, and there are $y-k\in(\frac{1}{2},\frac{3}{4}k)$ ones in the block of a corresponding solution of $y$. We pessimistically consider the decrease of $h$ by flipping $\frac{1}{4}k$ ones and keeping other bits unchanged of the corresponding solution. Hence, we have
$E[h$ ${}(P_{t})-h(P_{t+1})\mid P_{t},A_{3}]\geq\frac{1}{\mu}\frac{\binom{y-k}{k/4}}{n^{k/4}}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-k/4}\frac{1}{2}\geq\frac{1}{2e\mu}\frac{\binom{k/2}{k/4}}{n^{k/4}}.$
Therefore, we have
$E[h$ ${}(P_{t+1})-h(P_{t})\mid P_{t},v\notin P_{t}]=\sum_{i=1}^{3}\Pr[A_{i}]E[h(P_{t+1})-h(P_{t})\mid P_{t},A_{i}]$
[…]
where the third inequality uses $\binom{k/2}{k/4}\leq\binom{\frac{3}{4}k}{k/2}$ and $n^{k/2}\geq n^{k/4}$, and the last inequality uses $k\geq 1$. Since the maximum $h$ is $e\mu n^{k/2}+1$, we know that the Pareto front point $v$ will be reached in at most
$(e\mu n^{k/2}+1)\frac{{4e\mu n^{k/2}\sqrt{k}}}{4^{k/4}}=O\mathopen{}\mathclose{{}\left(\frac{\mu k^{1/2}}{2^{k/2}}\mu n^{k}}\right)$
in expectation, and will survive in all future populations. Consider the process of covering all non-inner Pareto front points in the order of covering the $1$-st level, …, $(m/2)$-th level. Since the 00-th level points are already covered from our assumption, and since there are at most $M$ non-inner Pareto front points, we know that the remaining levels are covered in at most $O(\frac{\mu k^{1/2}}{2^{k/2}}M\mu n^{k})$ iterations in expectation. With the same argument in the proof of Lemma 7, we know that the expected number of iterations to cover the full Pareto front is at most $eM\mu n^{k}$. Hence, we proved this lemma. ∎
@@
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with $\mu\geq 2(M+1)$ and stochastic population update to optimize the $m$OJZJ problem. Then after at most $\frac{1}{2}em^{k+1}k^{k}+en\mu M+O(\min\{\frac{\mu k^{1/2}}{2^{k/2}},1\}M\mu n^{k})$ iterations in expectation, the full Pareto front is covered.
Theorem 14 shows that the best runtime for the SMS-EMOA with stochastic population update is $O(\min\{\frac{Mk^{1/2}}{2^{k/2}},1\}M^{2}n^{k})$ when we choose $\mu=\Theta(M)$. Recall that the runtime of the original SMS-EMOA is $O(M^{2}n^{k})$ for $\mu=\Theta(M)$ from Theorem 8. We know the asymptotic speed-up is $\Theta(\max\{\frac{2^{k/2}}{Mk^{1/2}},1\})$ with $M=(2n/m-2k+3)^{m/2}$. Hence, when the objective number is not small, like at least the same size of $k$, then the speed-up is only $\Theta(1)$. We note that it is not contradictory to the nice speed-up obtained for two objectives in [BZLQ23] but adds a limitation situation for the stochastic population update.
6 Heavy-Tailed Mutation Helps
@@
6.1 Heavy-Tailed Mutation
Different from the standard bit-wise mutation, the heavy-tailed mutation flips each bit of the solution with probability $\alpha/n$ independently, where $\alpha$ follows a power-law distribution with parameter $\beta$ defined as follows.
@@ 6.1 Heavy-Tailed Mutation
Let $n\in\mathbb{N}$ and $\beta>1$. We say $\alpha$ follows the power-law distribution if for all $i\in[1..n/2]$, $\Pr[\alpha=i]=\mathopen{}\mathclose{{}\left(\sum_{j=1}^{n/2}k^{-\beta}}\right)^{-1}i^{-\beta}$.
This power-law distribution has a good feature that a point with Hamming distance of $j$ of a given bit-string will be generated with a probability of $\Omega(j^{-\beta})$. It will facilitate the jump with a larger size and facilitate escaping from the local optimum. The heavy-tailed mutation has resulted in an asymptotic performance speed-up by a factor of $k^{\Omega(k)}$ of the $(1+1)$ EA on the single-objective Jump function with gap size $k$ [DLMN17]. For the bi-objective OJZJ, the speed-up of $k^{\Omega(k)}$ is also witnessed when it is applied to the GSEMO [DZ21] and to the NSGA-II [DQ23a]. Several other positive theoretical results exist for this heavy-tailed mutation, or more generally, other heavy-tailed parameter choises [FQW18, WQT18, QGWF21, COY21, DELQ22, ABD22, DGI22, DR23]. Prior to this work, no theoretical analysis of the heavy-tailed mutation operator in many-objective optimization existed.
6.2 Runtime
@@
Since the heavy-tailed mutation does not change the survival selection of the original SMS-EMOA, we easily have the following corollary about the survival of the individuals with rank one from Lemma 4.

### 25 Jan 2024 v2

@@ Introduction
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with $\mu\geq M$ to optimize the $m$OJZJ problem. Then after at most $e\mu(mk)^{k}(1+\ln m)$ iterations in expectation, the population (and also the populations afterwards) contains at least one inner Pareto optimum.
In the next lemma, we consider the stage of covering all inner Pareto front points once at least one such point is in the population. The key to the proof is that as long as we have not yet discovered the full inner Pareto front, there always exists a missing Pareto front point that is a neighbor of a point that is already covered by the population. Hence choosing the right parent and flipping the right single bit suffices to cover the desired point on the Pareto front.
@@
The last stage is to cover the remaining Pareto front points. The following lemma shows the runtime. The key of the proof is that we can divide the Pareto front into several levels and any individual in the $i$-th level can be generated from a point in the $(i-1)$-th level by flipping the right $k$ bits.
@@
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with $\mu\geq M$ to optimize the $m$OJZJ problem. Then after at most $e\mu(mk)^{k}(1+\ln m)+e\mu Mn+e\mu Mn^{k}$ iterations ($\mu+e\mu(mk)^{k}(1+\ln m)+e\mu Mn+e\mu Mn^{k}$ function evaluations) in expectation, the full Pareto front is covered.
Since $k\leq n^{\prime}/2=n/m$ and $M\geq 2^{m/2}$, we easily see that both runtime expressions are $O(\mu Mn^{k})$, even when allowing $k$, $m$ and $\mu$ to depend on $n$.
Runtime of the GSEMO on $m$OJZJ
Since our arguments above can easily be extended to analyze the runtime of the GSEMO on the $m$OJZJ, we quickly do so for reasons of completeness. The GSEMO is a multiobjective counterpart of the ${(1+1)}$ EA. The initial population consists of a single randomly generated solution. In each iteration, a solution is uniformly at random picked from the population to generate an offspring via standard bit-wise mutation. If this offspring is not dominated by any solution in the population, it is added to the population and all solutions weakly dominated by it are removed. It is not difficult to see that all solutions in the population of the GSEMO are mutually non-dominated. Hence, the population size of the GSEMO is at most $M$ by Lemma 3.
When analyzing the runtime of the GSEMO on $m$OJZJ, the main difference is that SMS-EMOA requires a statement like Lemma 4 to ensure that previous progress is not lost via unlucky selection decisions. For the GSEMO, this property follows immediately from the selection mechanism, which keeps all non-dominated solutions. Together with the upper bound $M$ of the population size, we obtain the following theorem.
@@
Let $k\leq n^{\prime}/2$. Consider using the GSEMO to optimize the $m$OJZJ problem. Then after an expected number of at most $eM(mk)^{k}(1+\ln m)+eM^{2}n+eM^{2}n^{k}$ iterations (or $1+eM(mk)^{k}(1+\ln m)+eM^{2}n+eM^{2}n^{k}$ fitness evaluations), the full Pareto front is covered.
Reduced Impact of Stochastic Population Update
Bian, Zhou, Li, and Qian (2023) proposed a stochastic population update mechanism for the SMS-EMOA and proved that, somewhat unexpectedly given the state of the art, it can lead to significant performance gains. More precisely, it was proven that the classic SMS-EMOA with reasonable population size $\mu=\Theta(n)$ solves the bi-objective OJZJ problem in an expected number of $O(n^{k+2})\cap\Omega(n^{k+1})$ function evaluations, whereas for the SMS-EMOA with stochastic population update, $O(n^{k+2}\min\{1,2^{-k/4}n\})$ suffice. Hence, for $k=\omega(\log n)$, a super-polynomial speed-up was shown.
In this section, we extend the analysis of (Bian et al. 2023) to $m$ objectives. Unfortunately, we will observe that the larger population sizes necessary here reduce the impact of the stochastic population update. As in (Bian et al. 2023), we have no proven tight lower bounds for the SMS-EMOA with stochastic population update, but our upper-bound proofs suggest that the reduced effect of the stochastic population update on the runtime guarantee is real, that is, the impact on the true runtime is diminishing with the larger population sizes necessary in the many-objective setting.
Stochastic Population Update
The rough idea of the stochastic selection proposed by Bian et al. (2023) is that a random half of the individuals survive into the next generation regardless of their quality. The individual to be discarded is chosen from the remaining individuals according to non-dominated sorting and hypervolume contribution as in the classic SMS-EMOA. This approach resembles the random mixing of two acceptance operators in the Move Acceptance Hyper-Heuristic studied recently (Lehre and Özcan 2013; Lissovoi, Oliveto, and Warwicker 2023; Doerr et al. 2023).
More precisely, after generating the offspring $x^{\prime}$, the SMS-EMOA with stochastic selection chooses from $P_{t}\cup\{x^{\prime}\}$ a set $R^{\prime}$ of $\lfloor(\mu+1)/2\rfloor$ solutions randomly with replacement. The individual $z^{\prime}$ to be removed from $P_{t}\cup\{x^{\prime}\}$ is then determined via non-dominated sorting and hypervolume contribution applied to $R^{\prime}$ only. We note that with this mechanism, even the worst solution enters the next generation with probability at least $1/2$.
Runtime
To analyze the runtime of the SMS-EMOA with stochastic population update, we first derive and formulate separately two insights on the survival of solutions. The following Lemmas 10 is an elementary consequence of the algorithm definition, whereas Lemma 11 builds on Lemma 4.
@@
Consider using the SMS-EMOA with stochastic population update to optimize the $m$OJZJ problem. For any iteration $t$ and $x\in R_{t}$, we have that there is a $y$ with $f(y)=f(x)$ in $P_{t+1}$ with probability at least $1/2$.
@@
Consider using the SMS-EMOA with stochastic population update and with $\mu\geq 2(M+1)$ to optimize the $m$OJZJ problem. If at some time $t$ the combined parent and offspring population $R_{t}$ of the SMS-EMOA contains some solution $x$, then at any later time $s>t$ the parent population $P_{s}$ contains a solution $y$ such that $y\succeq x$.
Since the proofs of Lemmas 5 and 6 for the classic SMS-EMOA mostly relied on elementary properties of standard bit-wise mutation and on the survival guarantee of Lemma 4, we can now use the survival guarantee of Lemma 11 to obtain analogous results for the SMS-EMOA with stochastic selection (at the price of requiring essentially twice the population size). This yields the following estimates for the time to obtain at least one inner Pareto optimum and the time to cover all inner Pareto front points starting from a population with at least one inner Pareto optimum.
@@
Let $k\leq n^{\prime}/2$. Consider using the SMS-EMOA with stochastic population update and with $\mu\geq 2(M+1)$ to optimize the $m$OJZJ problem. Then
  • after at most $e\mu(mk)^{k}(1+\ln m)$ iterations in expectation, the population (and also the populations afterwards) contains at least one inner Pareto optimum;
  • after another at most $en\mu M$ iterations in expectation, all inner Pareto front points are covered.
Now we consider the runtime for the full coverage of the Pareto front after all inner Pareto front points are covered. Similar to the proof of Lemma 7 for the original SMS-EMOA, we partition the Pareto front points into $m/2+1$ levels based on the number of extreme blocks they contain. The difference here is that we consider the effect of the gap points, similar to Bian et al. (2023).
@@
Let $k\leq n^{\prime}/2$. Consider using the SMS-EMOA with stochastic population update and with $\mu\geq 2(M+1)$ to optimize the $m$OJZJ problem. Assume that the current population covers all inner Pareto front points. Then after at most $O(\min\{\frac{\mu k^{1/2}}{2^{k/2}},1\}M\mu n^{k})$ iterations in expectation, the full Pareto front is covered.
@@
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with $\mu\geq 2(M+1)$ and stochastic population update to optimize the $m$OJZJ problem. Then after at most $e\mu(mk)^{k}(1+\ln m)+en\mu M+O(\min\{\frac{\mu k^{1/2}}{2^{k/2}},1\}M\mu n^{k})$ iterations in expectation, the full Pareto front is covered.
Comparing the runtime guarantees or Theorem 8 (classic SMS-EMOA) and Theorem 14 (SMS-EMOA with stochastic selection), we see that stochastic selection can at most lead to a speed-up by a factor of order $2^{k/2}/\mu k^{1/2}$. Now $\mu$ is at least $\Omega(M)=\Omega((2n/m-2k+3)^{m/2})$. Consequently, the advantage of stochastic selection is rapidly decreasing with growing numbers of objectives and vanishes, e.g., when $m\geq k$.
Heavy-Tailed Mutation Helps
@@
Heavy-Tailed Mutation
Different from the standard bit-wise mutation operator, which flips each bit independently with probability $1/n$, the heavy-tailed mutation operator proposed in (Doerr et al. 2017) flips each bit independently with probability $\alpha/n$, where $\alpha$ follows a power-law distribution with parameter $\beta$. The number $\alpha$ is sampled anew in each application of the heavy-tailed mutation operator. The underlying power-law is defined as follows.
@@
Let $n\in\mathbb{N}$ and $\beta>1$. We say $\alpha$ follows a power-law distribution with (negative) exponent $\beta$ if for all $i\in[1..n/2]$, we have $\Pr[\alpha=i]=\left(\sum_{j=1}^{n/2}k^{-\beta}\right)^{-1}i^{-\beta}$.
The power-law feature extends to the heavy-tailed mutation operator in the sense that it generates an offspring with Hamming distance $j$ from the parent with probability $\Omega(j^{-\beta})$. This facilitates larger jumps in the search space, and thus escaping from a local optimum.
The heavy-tailed mutation operator has resulted in asymptotic performance gains by a factor of $k^{\Omega(k)}$ for the ${(1+1)}$ EA optimizing single-objective (classic and generalized) Jump functions with gap size $k$ (Doerr et al. 2017; Bambury, Bultel, and Doerr 2021). For the bi-objective OJZJ benchmark, again a speed-up of $k^{\Omega(k)}$ was proven when optimized via the GSEMO (Zheng and Doerr 2023d) and the NSGA-II (Doerr and Qu 2023a). Several other positive theoretical results exist for this heavy-tailed mutation, or more generally, other heavy-tailed parameter choices (Friedrich, Quinzan, and Wagner 2018; Wu, Qian, and Tang 2018; Quinzan et al. 2021; Corus, Oliveto, and Yazdani 2021; Dang et al. 2022; Antipov, Buzdalov, and Doerr 2022; Doerr, Ghannane, and Ibn Brahim 2022; Doerr and Rajabi 2023).
Prior to this work, no theoretical analysis of the heavy-tailed mutation operator in many-objective optimization existed.
Runtime
@@
Since the heavy-tailed mutation does not change the survival selection of the original SMS-EMOA, we immediately have the following result on the survival of the individuals.